# Anonymization by Design of Language Modeling

### v1

@@ 6.3. Utility and Privacy Tradeoff for Causal Language Modeling
This section evaluates the utility and privacy trade-off of our PPclm-gpt compared to other comparative baselines. Figure 12 presents for the N2c2 and the Hcl datasets the utility through the prism of token accuracy as a function of privacy assessed by the model’s ability to predict both direct and indirect identifiers during generation. The results show similar observations to those of PPmlm-bert but with much better privacy for PPclm-gpt compared to other baselines (up to 20% more privacy). Without any protection (i.e., clm-gpt which corresponds to a classic fine-tuning), the model learns the identifiers during training. After 64 epochs, this model is even able to regurgitate the direct and indirect identifiers it saw during training (revealed by a token accuracy close to 1). IPPclm-gptert (i.e., protecting only indirect identifiers) provides a similar level of privacy to classic fine-tuning because indirect identifiers are small in number. However, the utility is limited to 80% for the N2c2 dataset for instance, which means that the model does not regurgitate indirect identifiers because they were not learned during training. DPPclm-gpt and clmA-gpt, in turn, provide similar results with privacy decreasing to 80% privacy after 64 epochs. This decrease means that the model learns indirect identifiers over the epoch and tends to regurgitate them more frequently.
(a) N2C2
(b) HCL
@@
(a) Direct Privacy
(b) Indirect Privacy
Figure 13. Evaluating the utility and privacy tradeoff for CLM by biasing the privacy metric toward direct identifiers or indirect identifiers for the N2c2 dataset.
Figure13 illustrates the tradeoff between utility and privacy when privacy is measured solely through quantifying the leakage of direct (Figure13(a)) or indirect (Figure13(b)) identifiers. The results are similar to the observations made with masked language modeling. More precisely, when privacy only takes into account direct identifiers (i.e. this corresponds to pseudonymization), it is the baselines that protect direct identifiers that provide the best compromise. Conversely, if privacy only takes into account indirect identifiers, it is the baselines that protect indirect identifiers that provide the best compromise. This clearly shows that pseudonymization is not sufficient to protect a model against identifiable memorization, and that protection of both direct and indirect identifiers like done by our solution is necessary. Finally, the use of identifying words in the context for prediction has almost no impact on the utility and privacy tradeoff (comparison between DPPclm-gpt and clmA-gpt).
@@ 7. Limitations and discussions
The work presented in this paper represents a practical step towards protecting the training data against privacy leakage through the LLMs. There are, however, some limitations in several areas. Although the study and results have been validated on two datasets, these datasets are small in size. A validation on a larger dataset requires more computational resources and is left as future work.
@@
Another limitation is that the identification of direct identifiers of our privacy-preserving scheme relies on a NER, and the quality of this NER (Named Entity Recognition) can impact the performance of the protection of our solution. This limitation is common to all pseudonymizations relying on a NER and affects both the protection itself by missing identifying words, and the pseudonymization measure by not counting certain leaks of personal information. However, our solution is not impacted by misspelled and/or foreign words. Indeed, a common problem of NER in free text is the management of misspelled and/or foreign words that cannot be found in dictionaries. These words can be identifiers and not be protected. It is worth noting that the advantage of our solution is that these words will be identified as indirect identifiers in our solution and will not be memorized.
The preprocessing step of our solution to identify indirectly identifying words is somewhat like implementing k-anonymity with k=2. More precisely, this means that a term used by only one individual will be protected (i.e., not memorized by the model). That is, the words are at worst used in the reports of two patients and it is not possible for an attacker to easily identify either patient. We can easily change this value to a larger value. For example, by taking k=3, we ensure that the words that will be predicted by the model are at worst used in the reports of three patients, making the attacker’s task even more difficult. This configurable protection (i.e., the value of k) does not include any additional cost and only consists of navigating the bipartite graph built in the preprocessing step of our solution.
@@ 8. Conclusion
This article presents a methodology to avoid a model having memorized directly or indirectly identifying information linked to the specialization corpus used to specialize it. This absence of identifier memorization allows it to be shared without risk that the model can regurgitate re-identifying information during its exploitation. Our privacy preserving methodology of model specialization takes into account masked language modeling (i.e., for BERT-type models) and causal language modeling (i.e., for GPT-type models). We exhaustively experiment our approach and show that it offers the best compromise between utility and privacy.

### v2

@@ 6.3. Utility and Privacy Tradeoff for Causal Language Modeling
This section evaluates the utility and privacy trade-off of our PPclm-gpt compared to other comparative baselines. Figure 9 presents the utility through the prism of token accuracy as a function of privacy assessed by the model’s ability to predict both direct and indirect identifiers during generation. The results show similar observations to those of PPmlm-bert but with much better privacy for PPclm-gpt compared to other baselines (up to 20% more privacy). Without any protection (i.e., clm-gpt which corresponds to a classic fine-tuning), the model learns the identifiers during training. After 64 epochs, this model is even able to regurgitate the direct and indirect identifiers it saw during training (revealed by a token accuracy close to 1). IPPclm-gpt (i.e., protecting only indirect identifiers) provides a similar level of privacy to classic fine-tuning because indirect identifiers are small in number. However, the utility is limited to 80%, which means that the model does not regurgitate indirect identifiers because they were not learned during training. DPPclm-gpt and clmA-gpt, in turn, provide similar results with privacy decreasing to 80% privacy after 64 epochs. This decrease means that the model learns indirect identifiers over the epoch and tends to regurgitate them more frequently. Finally, DP drastically impacts the model’s utility. Results show that token accuracy is lower than $0.3$. Even worse, since DP is content-independent, privacy leakage increases slightly with the number of epochs, indicating that the model tends to increase memorization and regurgitation of identifiers as fine-tuning progresses. In Appendix A.2, we examine the behavior of PPclm-gpt and other baselines if privacy leakages are measured only from direct or indirect identifiers.
(a) BLEU score
(b) ROUGE score
Figure 10. A GPT model fine-tuned through our privacy-preserving language modeling provides roughly the same BLUE and ROUGE score than a model fine-tuned with speudonymized data, while a model without any protection provides better utility (15% after 32 epochs).
We now evaluate the usefulness of the fine-tuned GPT models through text generation quality assessment metrics (Figure 10). More precisely, we leverage the BLUE and the ROUGE scores which rely on n-grams overlap to measure similarity and recall between the text generated by the model and the reference text (i.e., the original one in our case). Figure 10(a) and Figure 10(b) depict these scores over the number of epochs for a fine-tuned GPT model without any protection (i.e., clm-gpt), for a GPT model fine-tuned with pseudonymized data (i.e., clmA-gpt), and for our privacy-preserving modeling scheme PPclm-gpt. Results show that for both utility metrics, clmA-gpt and PPclm-gpt provide roughly the same level of utility, while the solution without protection improves by 15% the utility after 32 epochs.
6.4. Impact of Membership Inference Attacks
To evaluate the accuracy of the membership inference attacks we conducted a ROC (Receiver Operating Characteristic) analysis which consists of comparing the positive predictions (i.e., membership guesses) over negative (non-membership) predictions. Specifically, we consider the true-positive rate (TPR) at low false-positive rates (FPR) for an attack conducted on a fine-tuned BERT model. The ROC curve (Figure 11) depicts the AUC (Area Under the Curve) for a model specialized without protection (i.e., mlm-bert), a model specialized from speudonimized data (i.e., mlmA-bert), and a model specialized following our scheme (i.e., PPmlm-bert). The dashed line shows AUC=0.5 which corresponds to random guess. Results show that for low false-positive rates (i.e., FPR at 1%), the true-positive rate is around 0.9% for both mlm-bert and PPmlm-bert, and around 2.5% for mlmA-bert. The AUC of PPmlm-bert is slightly lower than the AUC of both mlm-bert and mlmA-bert (i.e., 0.59 versus 0.61) that tends to say that the MIA applied on the model fine-tuned following our privacy-preserving language modeling scheme is slightly less efficient than on fine-tuned models without protection or using pseudonymized data.
Figure 11. ROC curve for the Membership Inference Attack with log-log scaling to emphasize the low-FPR regime.
A model can predict or regurgitate identifying terms. Let us now see if this prediction of identifiers can be used to infer membership at low cost, i.e., an individual will be considered as a member if at least one of their indirect identifiers is predicted by the model. Figure 12 depicts the precision, recall, and F1-Score of this low-cost membership inference when the adversary is able to construct the same bipartite graph as the one used in the fine-tuning (i.e., the adversary has the data used to fine-tune the model, which corresponds to a privacy audit carried out by the designer on his own model). The results for a fine-tuned model without protection show that it is possible to infer membership with high precision and a recall close to 0.7. Conversely, when the model is fine-tuned with our privacy-preserving solution, the recall drops to 0.1, drastically reducing the attacker’s ability to infer membership.
Figure 12. Inferring membership at low cost: by inspecting the identifiers predicted or regurgitated by the model, an adversary is able to predict membership with a good accuracy.
Figure  13, in turn, illustrates the F1-Score of the membership inference attack as a function of the number of epochs (i.e., overfitting) for unprotected language modeling (i.e., mlm-bert) and our privacy-preserving scheme (i.e., PPmlm-bert). The results show that, without protection, membership inference increases with the number of epochs, indicating a strong correlation between overfitting and the ability to predict membership. Conversely, membership inference for PPmlm-bert remains stable as training progresses.
Figure 13. While membership inference increases with the number of epochs (i.e., overfitting) in the case of unprotected language modeling, our privacy-preserving language modeling scheme is not affected by overfitting.
6.5. Considering n-grams
The proposed solution only considers isolated words as indirect identifiers. However, these indirect identifiers can be composed of multiple terms (i.e., n-grams of variable size n). Extending our solution to consider these n-grams as identifiers involves generating a bipartite graph accounting them. We first analyze the dataset to quantify the number of n-grams detected as indirect identifiers (i.e., unique to a patient). Figure 14 illustrates the distribution of the number of n-grams as indirect identifiers per patient for different values of n. Results show that, on average, the number of indirect identifiers doubles if we consider n from 1 (i.e., a word unique to a patient) to 2 (i.e., a sequence of two words unique to a patient). Considering n greater than 2 does not significantly increase the number of indirect identifiers.
Figure 14. Distribution of the number of indirect identifiers by considering n-grams as potential identifiers: the number of indirect identifiers does not increase after n=2.
(a) MLM
(b) CLM
Figure 15. By leveraging the bipartite graph to account the n-grams as potential indirect identifiers, our privacy-preserving language modeling can avoid the prediction or the regurgitation of these identifiers.
We then extend our solution to consider n-grams as potential indirect identifiers. Figure 15 quantifies the number of identifiers predicted or generated by the fine-tuned BERT or GPT models for n from 1 (i.e., a single word) to 3 (i.e., a unique sequence of 3 words). We plot the results for models accounting both n=1 (dash lines) and n=3 (continuous lines) during our language modeling, and models using comparative language modeling baselines. First results show that for BERT and GPT models fine-tuned without protection (i.e., mlm-bert or clm-gpt) or fine-tuned with pseudonimized data (i.e., mlmA-bert and clmA-gpt), the number of identifiers increases according to the value of n. For instance, for clm-gpt, the number of identifiers composed of a single word is around 30,000, while this number increases to more than 80,000 when the identifiers can also be composed of a sequence of two words. Similarly, when our language modeling scheme considered only single words as indirect identifiers during the fine-tuning, the model can still regurgitate sequences of two-word identifiers. However, when our language modeling scheme considered identifiers up to n=3 during the fine-tuning, the number of regurgitated identifiers remains stable, which is consistent because these 3-grams have been included in the blacklist of terms not to be memorized by the model.
@@ 7. Limitations and discussions
The work presented in this paper represents a practical step towards protecting the training data against privacy leakage through the fine-tuning of pre-trained LLMs. There are, however, some limitations in several areas. Although the study and results have been validated on one dataset, this dataset is small in size. A validation on a larger dataset requires more computational resources and is left as future work.
@@
Another limitation is that the identification of direct identifiers of our privacy-preserving scheme relies on a NER (Named Entity Recognition), and the quality of this NER can impact the performance of the protection of our solution. This limitation is common to all pseudonimization schemes relying on a NER and affects both the protection itself by missing identifying words, and the privacy metrics by not counting certain leaks of personal information. However, our solution is not impacted by misspelled and/or foreign words. Indeed, a common problem of NER in free text is the management of misspelled and/or foreign words that cannot be found in dictionaries. These words can be identifiers and not be protected. It is worth noting that the advantage of our solution is that these words will be identified as indirectly identifiers in our solution and will not be memorized.
It is also worth noting that the cost of our preprocessing step (i.e., building the bipartite graph) is significantly lower than using NER and fine-tuning the model. For example, when fine-tuning a BERT model (with 64 epochs) using a large dataset (50,000 documents), building the bipartite graph takes 0.3% of the overall fine-tuning time, while using NER takes 7%.
The language modeling scheme presented here avoids the model from memorizing direct and indirect identifying information present in the training data (including medical datasets and sensitive data) used for the specialization of off-the-shelf foundation LLMs. On the basis of the article 9 of the GDPR and on the Opinion 28/2024 of the EDPB on certain data protection aspects related to the processing of personal data in the context of AI models [^10] https://www.edpb.europa.eu/system/files/2024-12/edpb_opinion_202428_ai-models_en.pdf as well as the AI Privacy Risks & Mitigations Large Language Models (LLMs) report [^11] https://www.edpb.europa.eu/our-work-tools/our-documents/support-pool-experts-projects/ai-privacy-risks-mitigations-large_en, it is possible to say the fine-tuning of the LLM does not give birth to an fully anonymised model, due to the fact that the full anonymisation of the model can only be reached by an anonymisation process during the data preparation and preprocessing and the training phases of the foundation model [^12] AI Privacy Risks & Mitigations Large Language Models (LLMs) report, p. 25 et p. 33. Due to the fact that the model, before fine-tuning, was not, from the outset, anonymised, what is obtained is anonymization of the secondary learning data set used during fine-tuning, but not anonymization of the set of personal data on the basis of which the foundation model learned beforehand. In other words, a lot of personal data can be regurgitated by the fine-tuned model, even if it is not the health data on the basis of which the fine-tuning took place. Furthermore, interactions between the training data of the foundation model and the data used for fine-tuning are not controlled and may give rise to hallucinations. Consequently, the requirement set out in the EDPB opinion of an ”insignificant probability of personal data regurgitation” to qualify the fine-tuned model as anonymous is not met (Opinion 28/2024 on certain data protection aspects related to the processing of personal data in the context of AI models). Consequently, this means that the fine-tuned model will have to respect the conditions set out by the EDPB, in the three scenarios in the presence of a re-use described in their opinion  [^13] Opinion 28/2024 on certain data protection aspects related to the processing of personal data in the context of AI models, p. 4..
@@ 8. Conclusion
This article presents a methodology to prevent a pre-trained LLM from memorizing direct or indirect identifiers linked to individuals present in a dataset used to specialize it. This absence of identifier memorization facilitates model sharing by limiting the risk of it regurgitating identifying information during further exploitation. Our privacy-preserving language modeling scheme takes into account masked language modeling (i.e., for BERT-like models) and causal language modeling (i.e., for GPT-like models). We exhaustively evaluated our approach and demonstrated that it offers the best tradeoff between utility and privacy compared to other baselines.
Appendix A Privacy Leakage on direct or indirect identifiers
A.1. MLM
A model can leak information by predicting or regurgitating direct or indirect identifiers. Let us analyze the behavior of PPmlm-bert and the other baselines if we consider only the prediction of direct identifiers or indirect identifiers as privacy leakage (Figure 16). In the former case (Figure 16(a)), results show that mlm-bert and IPPmlm-bert exhibit a close utility and privacy tradeoff. Indeed, since indirect identifiers are not taken into account by the privacy measure, our mechanism for not memorizing indirect identifiers has no observable effect. Consequently, only the solutions that protect direct identifiers exhibit high privacy (i.e., PPmlm-bert, mlmA-bert, and DPPmlm-bert). However, we observe that mlmA-bert exhibits the lowest utility. This lower utility comes from the fact that all direct identifiers are pseudonymized in the training data and are not available in the context to improve the model’s prediction. Once again, PPmlm-bert offers the best trade-off between utility and confidentiality because it allows to benefit from direct identifiers in the context to improve predictions. By comparing the trade-off between mlmA-bert and DPPmlm-bert, we can evaluate the impact of memorization related to the use of directly identifying words in the context for mask prediction. We can see that the privacy impact is very small, and that using a context that exploits direct identifiers slightly improves the utility.
Similarly, if the privacy measure only captures the prediction of indirect identifiers as a privacy leakage (Figure 16(b)), only solutions that carefully protect this category of identifiers exhibit high privacy, this is the case for PPmlm-bert and IPPmlm-bert. The privacy degradation as a function of the number of epochs for the other baselines highlights the fact that they progressively learn indirect identifiers and predict them more and more often during training.
A.2. CLM
Figure17 illustrates the trade-off between utility and privacy when privacy is measured solely through quantifying the leakage of direct (Figure17(a)) or indirect (Figure17(b)) identifiers. The results are similar to the observations made with masked language modeling. More precisely, when privacy only takes into account direct identifiers (i.e., this corresponds to pseudonymization), it is the baselines that protect direct identifiers that provide the best compromise. Conversely, if privacy only takes into account indirect identifiers, it is the baselines that protect indirect identifiers that provide the best compromise. This clearly shows that pseudonymization is not enough to protect a model against identifiable memorization, and that protection of both direct and indirect identifiers like done by our solution is necessary. Finally, the use of identifying words in the context for prediction has almost no impact on the utility and privacy tradeoff (comparison between DPPclm-gpt and clmA-gpt).